A perfectly competitive firm produces 1000 units of burger in the long run. The marginal revenue is rm6. Calculate the firm's marginal cost, average fixed cost, long run average cost, total cost, total revenue and total profit
1
Expert's answer
2019-12-02T09:17:42-0500

In Perfectly Competitive market, Marginal revenue is equal to the price of the commodity. That means Price of the commodity is $6, which is equal to Marginal cost in case of perfect competition due to Marginal Cost pricing.

In perfect competition, Total Cost = MC*Q = 1000*6 = 6000

Average Fixed Cost = TFC/Q, as there is no fixed cost, burgers being the part of variable cost, AFC becomes zero.

Average run total cost = TC/Q = 6000/1000 = 6

Total Revenue, TR = P.Q = 6*1000 = 6000

Total Profit for perfect competition firm is

"\prod = TR - TC"

"\prod = 6000 - 6000 = 0"